OpenStudy (anonymous):

Find the area of an equilateral triangle (regular 3-gon) with the given measurement. 6-inch side. A = sq. in.

OpenStudy (anonymous):

All sides are 6 ins All angles are 60° A = (1/2) ( b ) ( h ) A = (1/2) ( b ) ( b sin 60° ) A = (1/2) b ² (√3 / 2) A = (1/4) √3 (36) A = 9√3 in²

OpenStudy (anonymous):

ty ^.^

OpenStudy (anonymous):

Anytime ♥